5x+8=3x-3 solve and state a property of equality
2(4y+3)=16 solve and state a property of equality
2/3p+1/5=4/5 solve and state a property of equality
3/4a-1/6=2/3a+1/4 solve and state a property if equality

Answer by addingup(3677)

5x+8=3x-3
2x=-11
x = -11/2
The subtraction property of equality states that subtracting the same number from each side of the equation gives us an equivalent equation.
-------------------------
You do the others, you need to learn and to learn you need to practice. The addition, multiplication, and division properties of equality all say the same: do to one side what you do to the other and you'll have an equivalent equation.
-----------------------
Other properties of equality:
Reflexive property, any quantity is equal to itself. e.g.: 16 = 4+4+4
Transitive property, says that if a = b and b = c then a = c